OpenStudy (anonymous):

What is one of the solutions to the following system of equations? y2 + x2 = 53 y − x = 5 (−10, −5) (−7, −2) (7, 2) (5, 10)

OpenStudy (campbell_st):

well I'd use substitution take the 2nd equation and make y the subject.. y = x + 5 and substitute it into the 1st equation \[x^2 + (x + 5)^2 = 53\] then you get after simplifying \[2x^2 + 10x -28 = 0\] divide each term by 2 and you have \[x^2 + 5x - 14 = 0\] just solve by factoring... and you'll get 2 solutions

OpenStudy (campbell_st):

but only 1 answer is shown in your choices.